A motor boat takes 4 hours to travel 352 km going upstream. The return trip takes 2 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?
Found 2 solutions by dewasmaharaja, nyc_function:
Answer by dewasmaharaja(5)

Let speed of boat in still water be x km/hr. and speed of stream be y km/hr.
Therefore,
Upstream, effective speed of boat = x - y = 352/4 = 88. This means, x = y + 88
Downstream, effective speed of boat = x + y = 352/2 = 176
Substituting value of x,
y + 88 + y = 176
2y = 176 - 88 = 88
Speed of stream, y = 44 km/hr
Speed of boat in still water, x = 44 + 88 = 132 km/hr

Answer by nyc_function(2741)

Let b = speed of boat in still water
Let r = speed of current
We use D = rt.
Upstream:
rate = b + r
time = 4
distance = 352
Downstream:
rate = b - r
time = 2
distance = 352
From this information we make two equations in two unknowns.
I will set up the equations and you can finish.
4(b + r) = 352...Equation A
2(b - r) = 352...Equation B
This is a good start for you. I'm sure you can finish now.